Find out the inductor current equation for an RL circuit with step response, assume the initial current through the inductor is 𝐼0.


1
Expert's answer
2022-03-13T18:49:49-0400

Circuit equation


"0=iR+L\frac{di}{dt}"

Solution to this equation


"i=Ae^{-t/\tau}"

Where "A - constant, \tau = L/R"

Initial conditions t=0; i=I0

Then "A=I_0"

Answer


"i=I_0e^{-\frac{t}{\tau}}"